Write a MATLab code for a function m-file that that can take three inputs (f, y, t) and uses a while loop to examine f(x) for x = y, x = y+2t, x = y+4t, ... until a number is larger than the previous value and returns the x value for which this happens.

Notes: Assume that f is decreasing at x=y but that there is a cutoff x=b > y after which f starts increasing. Assume the input for f will be a function handle and that the inputs for y and t will be real numbers. The MATLab code must be in a function m-file (i.e. include the "function" command in the first line).
Answer #1

function xx = func_xyt(f,y,t)
i=4;
y1=f(y);%y1 will hold previous run
y2=f(y+2*t);%y2 holds current run
while y1>=y2%while previous run is more than current one
    y1=y2;%now, previous one holds the cuurent one
    y2=f(y+t*i);%current one calculated with new input
    i=2*i;%counter doubled
end
xx=y+(i/2)*t;
end